OpenStudy (anonymous):

For Rational Functions, How do I simplify and state the excluded values? here's my equation: 72m^2/48m^3

OpenStudy (anonymous):

I think your answer would be 3 ____ 2m

OpenStudy (anonymous):

Im not really sure

OpenStudy (anonymous):

how did u solve that?

OpenStudy (angela210793):

72/48=3/2 m^2/m^3=1/m so 3/2m

OpenStudy (anonymous):

I simplified 72/48 and eventually got 3/2. Then i divided m^2 and m^3 and got m^-1. Ans when you have a negative exponent it goes to the bottom so therefore i got 3/2m

OpenStudy (anonymous):

and*